OpenStudy (dls):

100 ml of N/10 NaOH is mixed with 100ml of N/5 HCL solution and the whole solution is made up to 1 L What is the pH of resulting solution

OpenStudy (dls):

@yrelhan4

OpenStudy (yrelhan4):

ionic equilibrium not my thing. i'll still try. give me some time.

OpenStudy (dls):

(Y)

OpenStudy (dls):

Number of moles of NaOH=0.01 Number of moles of HCl=0.02

OpenStudy (dls):

Concentration of NaOH=0.01 Concentration of HCl=0.02

OpenStudy (shubhamsrg):

@shubhamsrg

OpenStudy (dls):

urself?

OpenStudy (shubhamsrg):

just bookmarking myself !

OpenStudy (dls):

:D

OpenStudy (shubhamsrg):

moles of HCl left after reacting = 0.01 volume now = 1 L molarity = 0.01 /1 = 10^(-2) Ph = -log(10^-2) am getting 2. But its wrong In case someone ( @yrelhan4 ) can catch the flaw?

OpenStudy (dls):

nooooooooooo!

OpenStudy (dls):

i think I got it

OpenStudy (shubhamsrg):

oye oye! tell me

OpenStudy (dls):

hold on confirming smt

OpenStudy (yrelhan4):

looks good. :P

OpenStudy (dls):

lol @shubhamsrg why on earth 2 medals!

OpenStudy (yrelhan4):

thank you for the solution @shubhamsrg

OpenStudy (shubhamsrg):

There is only one reason! I AM @shubhamsrg ! B|

OpenStudy (dls):

tell me jitna maine likha sahi tha?

OpenStudy (shubhamsrg):